# Pranav Suryadevara Morris Plains, NJ, USA

### v1

@@ Abstract
The growth of machine learning (ML) workloads has underscored the importance of efficient memory hierarchies to address bandwidth, latency, and scalability challenges. HERMES focuses on optimizing memory subsystems for RISC-V architectures to meet the computational needs of ML models such as CNNs, RNNs, and Transformers. This project explores state-of-the-art techniques such as advanced prefetching, tensor-aware caching, and hybrid memory models. The cornerstone of HERMES is the integration of shared L3 caches with fine-grained coherence protocols and specialized pathways to deep learning accelerators like Gemmini. Simulation tools like Gem5 and DRAMSim2 are used to evaluate baseline performance and scalability under representative ML workloads. The findings of this study highlight the design choices and anticipated challenges, paving the way for low-latency scalable memory operations for ML applications.
@@ I Introduction
Machine learning (ML) models have become integral to modern computing, with applications spanning computer vision, natural language processing, and autonomous systems. These models, such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and Transformers, often require extensive computational resources and memory bandwidth. Consequently, optimizing memory hierarchies for ML workloads is critical to improving performance and efficiency [1, 2].
RISC-V, an open source Instruction Set Architecture (ISA), is gaining traction due to its flexibility and extensibility [3, 4]. However, the performance of RISC-V processors for ML workloads is hindered by memory bottlenecks, including latency, limited bandwidth, and lack of scalability [6]. The HERMES project aims to design a high-performance memory subsystem tailored to ML workloads on RISC-V architectures. Using techniques such as advanced prefetching [8], tensor-aware caching [9], and hybrid memory models [7], HERMES seeks to overcome these challenges and enhance the efficiency of ML computations.
This report presents a foundational study of HERMES, outlining the design considerations, initial simulation results, and challenges in achieving a scalable, low-latency memory hierarchy for ML workloads and evaluate the performance using simulation tools such Gem5 [10] and DRAWSim2[11].
@@ II-A Memory Bottlenecks in ML Workloads
ML workloads, especially deep learning models, are characterized by large datasets and frequent memory access patterns. Studies have identified key challenges that impact the efficiency of memory subsystems:
@@
Tensor operations, such as matrix multiplications, require frequent access to large datasets stored in memory. The latency involved in fetching data from DRAM can significantly degrade performance if not mitigated by efficient caching and prefetching mechanisms [1, 2, 5].
@@
ML models, particularly those involving convolutions and recurrent operations, demand high memory bandwidth. The concurrent fetching of weights, activations, and gradients often saturates the available bandwidth, leading to bottlenecks in data transfer. This issue is exacerbated in systems with limited on-chip memory [6, 7, 12].
@@
As ML models grow in size and complexity, memory subsystems must scale to accommodate increasing data loads. Current architectures struggle to maintain performance as the dataset size exceeds cache capacity, requiring frequent access to slower off-chip memory [10, 15].
Addressing these bottlenecks requires innovative memory hierarchies that can balance latency, bandwidth, and scalability for different ML workloads.
@@ II-B Existing Memory Solutions for ML
Advanced prefetching techniques predict data access patterns and load data into the cache before it is needed. Techniques such as Stride Prefetching and Machine Learning-Based Prefetching have proven effective in reducing latency for sequential and non-sequential data access patterns [8, 5]. For instance, Intel’s microprocessors incorporate sophisticated hardware prefetchers that dynamically adapt to workload behavior [9, 12].
@@
Combining DRAM with High-Bandwidth Memory (HBM) offers a balance between capacity and speed. HBM provides significantly higher bandwidth compared to traditional DRAM, making it suitable for ML workloads with high data transfer requirements [7, 16]. For example, NVIDIA’s GPUs use HBM to deliver the necessary bandwidth for training large neural networks [13].
These techniques have shown promise in improving memory performance for ML workloads, but their integration in RISC-V architectures remains limited.
@@ II-C RISC-V-Based Accelerators
The open source nature of RISC-V has led to the development of several accelerators designed for ML workloads. These accelerators benefit from custom memory pathways and specialized hardware support for tensor computations.
@@
Gemmini is a configurable deep learning accelerator for RISC-V systems. It provides efficient support for matrix multiplication and convolution operations, which are fundamental to ML workloads [6, 14]. The Gemmini design allows integration with custom memory pathways, facilitating efficient data transfer between caches and the accelerator.
@@ II-D Research Gaps
    Providing a scalable and efficient solution for deep learning models on RISC-V platforms.
@@ IV Experimental Setup
  • Bandwidth (GB/s): The rate at which data is transferred between memory and the processor/accelerator. Higher bandwidth indicates better handling of large data transfers.
@@
  • L1 Cache: 32 KB per core, 8-way set associative.
  • L2 Cache: 256 KB per core, 8-way set associative.
  • Shared L3 Cache: 8 MB, 16-way set associative.
@@
  • Simulation Tools: Gem5 for CPU and cache simulation, DRAMSim2 for DRAM modeling.
The configurations are designed to balance accuracy and simulation runtime while providing insights into the performance benefits of the HERMES memory hierarchy.
@@
This section presents the performance evaluation of the HERMES memory hierarchy using simulations. The key metrics analyzed include latency, bandwidth, cache hit rates, and energy consumption. The results are compared against a baseline RISC-V system to highlight the improvements offered by HERMES.
@@
The latency and bandwidth of the HERMES architecture were evaluated and compared to a baseline RISC-V system. The results are summarized in Table I.
@@
Figure 2 shows a significant reduction in latency with the HERMES configurations, while Figure 3 highlights the bandwidth improvements due to the hybrid memory model.
@@
Cache hit rates were evaluated to determine the effectiveness of tensor-aware caching in HERMES. Table II shows the results.
@@ Discussion
These improvements highlight the effectiveness of integrating advanced memory design techniques in RISC-V architectures to meet the demands of modern ML workloads.
@@ Acknowledgment
The author expresses sincere gratitude to Professor Ray Simar for his invaluable guidance, support, and encouragement throughout the development of this project. This work was completed as part of COMP 554: Computer Systems Architecture at Rice University. Professor Simar’s deep expertise in computer architecture and his commitment to student learning have been instrumental in shaping this project. His insights and feedback have significantly improved our understanding of memory systems and their optimization for machine learning workloads.

### v2

# HERMES: High-Performance RISC-V Memory Hierarchy for ML Workloads

@@ Abstract
The growth of machine learning (ML) workloads has underscored the importance of efficient memory hierarchies to address bandwidth, latency, and scalability challenges. HERMES focuses on optimizing memory subsystems for RISC-V architectures to meet the computational needs of ML models such as CNNs, RNNs, and Transformers. This project explores state-of-the-art techniques such as advanced prefetching, tensor-aware caching, and hybrid memory models. The cornerstone of HERMES is the integration of shared L3 caches with fine-grained coherence protocols equipped with specialized pathways to deep-learning accelerators such as Gemmini. Simulation tools like Gem5 and DRAMSim2 were used to evaluate baseline performance and scalability under representative ML workloads. The findings of this study highlight the design choices, and the anticipated challenges, paving the way for low-latency scalable memory operations for ML applications.
@@ I Introduction
Machine learning (ML) models have become integral to modern computing, powering applications across fields such as Computer Vision, Natural Language Processing, and Autonomous Systems. With these advancements, models such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Transformers, Large Language Models (LLMs), and Large Vision-Language Models (LVLMs) often require extensive computational resources and memory bandwidth. Hence as a result optimizing memory hierarchies for ML workloads is of significant importance for enhancing performance and efficiency [1, 2].
RISC-V, an open-source Instruction Set Architecture (ISA) initially developed at the University of California, Berkeley, is gaining traction due to its flexibility and extensibility [3, 4]. However, the performance of RISC-V processors for ML workloads is currently limited by memory bottlenecks, including latency, limited bandwidth, and lack of scalability [6]. The HERMES project aims to design a high-performance memory subsystem that is tailored to optimize ML workloads on RISC-V architectures. Using techniques such as advanced prefetching [8], tensor-aware caching [9], and hybrid memory models [7], HERMES seeks to overcome these challenges and enhance the efficiency of ML computations.
This report presents a foundational study of HERMES, outlining the design considerations involved, the initial simulation results obtained, and the challenges anticipated to achieve a scalable, low-latency memory hierarchy system for ML workloads. The project evaluates the performance of the system using simulation tools such as Gem5 [10] and DRAWSim2[11].
@@ II-A Memory Bottlenecks in ML Workloads
ML workloads, especially the deep-learning models, are characterized by large datasets and frequent memory access patterns. Studies have identified that the key challenges that impact the efficiency of memory subsystems are:
@@
Tensor operations, such as matrix multiplication, require frequent access to large datasets stored in memory. The latency factor involves fetching in data from DRAM which can significantly degrade performance if it isn’t mitigated by efficient caching and prefetching mechanisms [1, 2, 5].
@@
ML models, particularly those involving convolutions and recurrent operations, demand a high memory bandwidth. The concurrent fetching of weights, activations, and gradients often saturates the available bandwidth, leading to bottlenecks in the data transfer. This issue is exacerbated in systems with limited memory available on the chip [6, 7, 12].
@@
As ML models grow in size and complexity, memory subsystems must scale to accommodate the increasing data loads. Current architectures struggle to maintain performance as the size of the dataset exceeds the cache capacity, thus requiring frequent access to the slower off-chip memory [10, 15].
Addressing these bottlenecks requires innovative memory hierarchies that can balance latency, bandwidth, and scalability for a variety of ML workloads.
@@ II-B Existing Memory Solutions for ML
Advanced prefetching techniques predict data access patterns and load data into the cache before it is needed. Techniques such as Stride Prefetching and Machine Learning-Based Prefetching have been proven effective in reducing latency for sequential and non-sequential data access patterns [8, 5]. For instance, Intel’s microprocessors incorporate sophisticated hardware prefetchers that dynamically adapt to the workload behavior at hand [9, 12].
@@
Combining DRAM with High-Bandwidth Memory (HBM) offers a balance between capacity and speed. HBM provides significantly higher bandwidth compared to the traditional DRAM, making it suitable for ML workloads with high data transfer requirements [7, 16]. For example, NVIDIA’s GPUs use HBM to deliver the necessary bandwidth for training large neural networks [13].
These techniques have shown promise in improving memory performance for ML workloads, but their integration in a RISC-V architecture remains limited.
@@ II-C RISC-V-Based Accelerators
The open-source nature of RISC-V has led to the development of several accelerators designed for ML workloads. These accelerators benefit from custom memory pathways and specialized hardware support for tensor computations.
@@
Gemmini is a configurable deep-learning accelerator for RISC-V systems. It provides efficient support for matrix multiplication and convolution operations, which are fundamental to ML workloads [6, 14]. The Gemmini design allows integration with custom memory pathways, facilitating efficient data transfer between caches and the accelerator.
@@ II-D Research Gaps
    Providing a scalable and efficient solution for deep-learning models on RISC-V platforms.
@@ IV Experimental Setup
  • Bandwidth (GB/s): The rate at which data are transferred between the memory and the processor/accelerator. Higher bandwidth indicates better handling of large data transfers.
@@
  • L1 Cache: 32 KB per core, 8-way associative set.
  • L2 Cache: 256 KB per core, 8-way associative set.
  • Shared L3 Cache: 8 MB, 16-way associative set.
@@
  • Simulation Tools: Gem5 for CPU and cache simulation, and DRAMSim2 for DRAM modeling.
The configurations are designed to balance accuracy and simulation runtime while providing insight into the performance benefits of the HERMES memory hierarchy.
@@
This section presents a performance evaluation of the HERMES memory hierarchy using simulations. The key metrics analyzed include latency, bandwidth, cache hits,energy consumption. The results are compared with a baseline RISC-V system to highlight the improvements offered by HERMES.
@@
The latency and bandwidth of the HERMES architecture were evaluated and compared with a baseline RISC-V system. The results are summarized in Table I.
@@
Figure 2 shows a significant reduction in latency with the HERMES configurations, while Figure 3 highlights bandwidth improvements due to the hybrid memory model.
@@
The cache hit rates were evaluated to determine the effectiveness of tensor-aware caching in HERMES. Table II shows the results.
@@ Discussion
These improvements highlight the effectiveness of integrating advanced memory design techniques into RISC-V architectures to meet the demands of modern ML workloads.
@@ Acknowledgment
The author expresses sincere gratitude to Professor Ray Simar for his invaluable guidance, support, and encouragement throughout the development of this project. This work was completed as part of COMP 554: Computer Systems Architecture at Rice University. Professor Simar’s deep expertise in computer architecture and his commitment to student learning have been instrumental in shaping this project. His insights and feedback have significantly improved my understanding of memory systems and their optimization for machine learning workloads.